OpenStudy (anonymous):

Write a system of two equations using x and y for the following problem: The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (nottim):

You know to use variables?

OpenStudy (anonymous):

ummm....yea

OpenStudy (nottim):

ok. im pretty you know how to do this then. what have you done so far?

OpenStudy (anonymous):

A + B = 170 A - B = 26

OpenStudy (nottim):

yup

OpenStudy (anonymous):

okay so then wat do i do next?

OpenStudy (nottim):

now, rearrange one of the equations to solve for either a or b.

OpenStudy (nottim):

ok?

OpenStudy (anonymous):

okay so then will it be: 26+b=170?

OpenStudy (anonymous):

idk how tor do this ;(

OpenStudy (nottim):

no...i dont think so...

OpenStudy (nottim):

take A + B = 170. can you isolate a for me?

OpenStudy (anonymous):

umm...-b+170..idk

OpenStudy (nottim):

a=-b+170. good. now sub that into A - B = 26 (for a)

OpenStudy (anonymous):

So then would b=144

OpenStudy (nottim):

i dunno. maybe?

OpenStudy (anonymous):

wat do u mean maybe???

OpenStudy (nottim):

i didnt do the math myself

OpenStudy (nottim):

you have to divide that by 2...

OpenStudy (nottim):

-2.

OpenStudy (nottim):

and originally, it would have been -144, not 144

hero (hero):

You can do all of that or you can simply add both equations together: A + B = 170 A - B = 26 ___________ 2A = 192 A = 192/2 A = 96 B = A - 26 B = 96 - 26 B = 70

OpenStudy (anonymous):

so it would be 72?

hero (hero):

Well I obviously added wrong

OpenStudy (anonymous):

yah think lol ;)

hero (hero):

A + B = 170 A - B = 26 ___________ 2A = 196 A = 196/2 A = 98 B = A - 26 B = 98- 26 B = 72

OpenStudy (anonymous):

Okay thnk u @hero but @NotTim cna u still help me solve the second part?

hero (hero):

What second part?

hero (hero):

You were only supposed to find the two numbers. A is one number. B is the other number. A = 98 B = 72 Is there something else you need to find besides that?

OpenStudy (anonymous):

No i need to know how to find 98 word for word

hero (hero):

When solving systems of equations, there are five methods you can use: 1. Substitution 2. Elimination 3. Addition 4. Graphing 5. Slope-Intercept

hero (hero):

Which method are you confused about?

OpenStudy (anonymous):

okay this is wat i'm tlking about: how did u get 2A=196?

hero (hero):

I added the two equations together.

hero (hero):

The elimination method involves adding equations together to eliminate variables.

hero (hero):

In this case I added A + B = 170 to A - B = 26. When you add two equations together, you add like terms. Therefore: A + A = 2A B - B = 0 170 + 26 = 196 Thus 2A + 0 = 196 or simply 2A = 196

hero (hero):